# Decoding Chess Puzzle Play and Standard Cognitive Tasks for BCI: A Low-Cost EEG Study

### v1

@@ 3.5 Participant Information and Chess Player Skill
Chess participants also filled out a form indicating the frequency of their Chess play. The distribution of responses is shown in Figure 6 - a substantial majority (75%) of participants responded that they play at least once per week. After gathering our data, we assessed skill distribution of Chess players by examining the maximum puzzle difficulty each participant successfully completed, as illustrated in Figure 7. For reference, puzzles rated at 1400 Elo are the most frequently attempted puzzles in the lichess.org database (Lichess.org, 2022). While participant skill level could theoretically influence neural responses during puzzle solving, our preliminary analyses found no significant effects. Given these results, subsequent analyses do not investigate skill-based effects. Future research examining skill-based neural responses may benefit from recruiting a larger participant pool with broader representation across skill levels.
Figure 6: Frequency of Chess play across all Chess participants (N=17). Most players (75%) played Chess at least once per week.
Figure 7: Distribution of maximum Chess puzzle difficulty level achieved across participants, measured in Elo Glicko2 rating (see Section 3.3). Preliminary quantile-based analyses revealed no significant associations between maximum Elo rating and EEG response patterns (analyses not reported in main results).
3.6 EEG Data Collection and Preprocessing
@@ 3.6.1 Epoching the Data
3.6.2 Frequency Domain Transformation and Prefrontal Averaging
Each epoch was transformed into the frequency domain with mne’s Multitaper library (Percival and Walden, 1993; Candy, 2019). Following the work of So et al. (2017), we delineated frequency bands as follows: $\theta$ (4, 8), $\alpha_{1}$ (8, 11), $\alpha_{2}$ (11, 14), $\beta_{1}$ (14, 25), $\beta_{2}$ (25, 35), $\gamma_{1}$ (35, 40), and $\gamma_{2}$ (40, 45)555 While we recognize that these are not necessarily the most common waveband delineations used, they represent a similar paradigm within this work done in (So et al., 2017).. The Simpson integral (Fernández et al., 2021) was used to extract single values representing the total power in each frequency band. Given our focus on overall pre-frontal activation patterns, we averaged the frequency-domain data from both probes into a single value per frequency bin, optimizing signal-to-noise ratio while maintaining measurement validity for our research objectives. Resulting power values were log-transformed.
3.6.3 Data Collection Challenges and Solutions
Throughout our study we experienced connectivity issues with the MUSE device. In our initial implementation of the experimental protocol, this would cause complete data loss for a participant. To mitigate data loss, we implemented a reconnection protocol prior to the start of each task block (e.g., one block of N-Back, Stroop, mental workload, or Chess puzzles). This approach significantly reduced data loss overall, but some blocks were still lost for some participants. Table 1 presents detailed information on the number of captured trials and epochs captured per-participant per-task after data loss was taken into account.
Table 1: Detailed dataset sizes after exclusions and filtering. Data here are reported from both phases I and phase II. Features per Task, are the number of participants and (Num. P) mean Blocks per Participant (B/P). For each level of Workload (W-Load) within Task, the features are: mean Trials per Block per Participant (T/B/P), total number of Trials (T), mean Epochs per Block per Participant (E/B/P), and total Epochs (E). Note that, although the number of trials is similar across workload levels, the total number of epochs increases as workload increases, because there will be more 1 second samples in higher workload conditions. For statistical modeling we average all epochs for each participant at the block level (and within levels of workload for within-task workload classification); for machine learning, epochs within each block, each of which contain all workload levels, are grouped together and used exclusively in the training or testing set.
Task Num. P Mean B/P W-Load Mean T/B/P Total T Mean E/B/P Total E
Chess 0 5.7 556 45.3 3109
1 6.4 631 68.2 4690
16 5.4 2 10.2 989 129.6 8981
3 8.3 758 146.9 9582
Total 30.65 2934 389.86 26362
N-Back 0 22.3 1397 27.0 1058
1 22.4 1345 29.0 1050
11 4 2 23.1 1480 33.7 1328
3 90.9 1438 36.4 1332
Total 22.7 5660 126.1 4768
Rotation 0 5.8 388 11.1 433
1 6.1 411 14.2 579
11 3.9 2 6.2 409 17.2 717
3 5.8 387 16.7 715
Total 23.9 1595 59.2 2444
Stroop 0 33.9 2260 31.9 1302
10 3.9 1 35.2 2317 36.6 1391
Total 69.1 4577 68.6 2793
@@ 3.7 Workload Labels
3.8 Statistical Methodology
The initial part of our analysis employs statistical methods to examine the measured neurophysiological data for both within-task workload and between-tasks. Except where otherwise stated, we utilize Linear Mixed-Effects models to achieve this aim. In each model, we incorporate Participant ID and Task Block per-participant as grouping factors for random effects. We implemented the models with R’s lmerTest package (Kuznetsova et al., 2017), and validated them through diagnostic checks that included verification of approximate normality of residuals with Quantile-Quantile (Q-Q) plots, posterior predictive checks for model fit, and examination of homoscedasticity and linearity against fitted values using the performance package (Lüdecke et al., 2021). To quantify the magnitude of the fixed effects, we calculate partial eta-squared ($\eta_{p}^{2}$) values using the effectsize library (Ben-Shachar et al., 2020). Estimated marginal means and post hoc pairwise comparisons were conducted with the emmeans library (Lenth, 2024). We used the conventional threshold of $\alpha=0.05$ for all statistical tests in our analysis.
3.9 Modeling of Reaction Time and Correctness
To investigate behavioral metrics we modeled reaction time as a function of difficulty for all of the trials over all participants. For the N-Back, Stroop, and Rotation tasks, time is defined as the reaction time taken to press a key following the stimulus presentation. For the Chess puzzle task, we define time as the time required to solve the puzzle. Reaction time data was log-transformed for statistical modeling (Lo and Andrews, 2015). Omnibus test p-values for per-task models were corrected with Benjamini-Hochberg correction, as were all post-hoc comparisons. Prior to correction, pairwise comparisons were Tukey-adjusted to control the family-wise error rate. We also modeled correctness as a function of difficulty. Because correctness of a given trial is binary, we employed a Generalized Linear Mixed-Effect Model (glmer) with a binomial family function for the analysis; random effects structure is the same as described in Section 3.8; p-values were adjusted using the same method as for the Reaction Time tests. Data visualizations are created using the seaborn library (Waskom, 2021), and display group means accompanied by their 95% confidence intervals, which were calculated using bootstrapping methods that account for the hierarchical structure of the data.
3.10 Statistical Modeling of EEG Data
@@
3.11 Machine Learning Analyses
@@
3.11.1 Grouped and Subsampled Monte Carlo Cross-Validation
Due to data loss from the MUSE device, the fact that the Chess task was collected for as long as the other three tasks combined, and due to varying time-spans of the other three tasks, our dataset contains imbalanced samples both within and across participants. However, this imbalance is not representative of the brain data itself; it stems solely from the experimental paradigm. To develop an online algorithm for differentiating between brain-states or workload levels in practice, we would modify the protocol to create a balanced dataset. Therefore, to test our data with machine learning in the fairest way possible we employed a grouped and subsampled Monte Carlo Cross-Validation methodology for both the within-task workload machine learning models and for the between-task model. The data preparation pipelines are shown in detail in Figures 9 and 10. This method involves 1,000 iterations for the cross-task workload model, and 1,000666For all of our models, after conducting 1,000 iterations of Monte Carlo Cross-Validation we determined that the 95 confidence interval of the mean fell within ±0.5 percentage point., where each iteration follows these steps:
@@ 3.11.1 Grouped and Subsampled Monte Carlo Cross-Validation
    Initialize a Random Forest classifier using the ensemble.RandomForest function from scikit-learn (Pedregosa et al., 2011) on the training data using default parameters  [^7] Model configuration included 100 trees (n__\__estimators), unlimited maximum depth with nodes expanded until reaching pure leaf or minimum split threshold, minimum samples for split of 2, minimum samples per leaf of 1, and Gini criterion for measuring split quality. We opted for default parameters to establish a robust baseline implementation while avoiding potential overfitting from extensive parameter tuning; record both overall results as well as results of per-participant classification.
3.11.2 Within-Task Workload Machine Learning
@@
For within-task workload we model each task separately. Because in this case we do not need to classify across tasks, we can leverage the entire dataset, including all valid data from participants in both phases I and II. See Table 2 for detailed information on average train and test set sizes across the Monte Carlo iterations.
Table 2: Number of participants per-model, and average train and test set sizes (rounded down) over all Monte Carlo iterations of within-task workload machine learning analysis. Due to the data splitting process discussed in 3.11.1, for both test and train sets the labels in the model are equally balanced for each participant. Participants are likewise equally represented in both train and test sets.
Trial type Number of participants Average train set sample size Average test set sample size
Chess 17 3,983 1987
N-Back 11 1,964 622
Rotation 10 842 220
Stroop 10 1,476 404
3.11.3 Cross-Task Machine Learning
@@
For cross-task analysis we require data for each participant from all tasks, therefore we use only the data collected in phase II for this machine learning set. For this machine learning set, we had a total of 6 participants. Across all Monte Carlo iterations we had an average of 2,723 training samples and 776 testing samples. Due to the data splitting process discussed in 3.11.1, the training and testing samples were each evenly balanced within each participant for samples per label, and were likewise balanced for samples per-participant.
Figure 9: Visualization of one iteration of the data splitting steps for the Monte Carlo pipeline for within-task workload classification. Blocks of grouped trials for each participant are split into training and testing sets randomly, with 80% of blocks in training and the remaining 20% of blocks in testing. Trials within training and testing blocks are subsampled randomly to the lowest frequency of workload label. Training data are scaled, and $\mu$ and $\sigma$ from the training data are used to scale the test data for that participant. All participants’ training sets are then combined and subsampled to the lowest frequency participant identifier; likewise happens for the test sets. This pipeline is done separately for each Monte Carlo simulation for each trial type.
Figure 10: Visualization of one iteration of the data splitting steps for the Monte Carlo pipeline for cross-task classification. Blocks of grouped trials within each trial type for each participant are split into training and testing blocks randomly, with 80% of blocks in training and the remaining 20% of blocks in testing. All such training and testing blocks are combined for a given participant, and trials within the training and testing blocks are then subsampled (separately) to the lowest frequency label, and scaled. All such training and testing blocks for all participants and trial types are combined into single train and test sets, which are separately subsampled to the lowest frequency participant ID. F1 results are reported from both the precision and recall scores on the test set overall, and per-participant.
scikit-learn (Pedregosa et al., 2011) was used to implement the grouping strategy, Random Forest model, and correctness statistics. imblearn (Lemaître et al., 2017) was used for subsampling of the data.
We report the averages over all 1,000 models of the precision, recall, and macro F1-score.
4 Results
4.1 Reaction Time and Correctness
Figure 11: Performance and reaction time as a function of workload level for each of the four tasks. In general, as workload increases, correctness decreases and reaction time increases. Note that the temporal (right-side) scale y-scales for the three cognitive neuroscience tasks are the same (0-4000ms), however the same axis for the Chess task is from (6000-17000ms).
@@ 4.1 Reaction Time and Correctness
See Figure 11 for a visualization of reaction time and correctness data. Regarding temporal statistics, for all tasks, time to solve was significant: N-Back ($F_{3,6166.49}=795.91$, $p<0.001$, $\eta^{2}_{p}=0.28$) [^8] All p-values reported inline reflect adjustments for multiple comparisons. Tables present both unadjusted (p) and adjusted (p.adj) values, Chess ($F_{3,2906.35}=300.37$, $p<0.001$, $\eta^{2}_{p}=0.24$), Rotation ($F_{3,1554.14}=127.65$, $p<0.001$, $\eta^{2}_{p}=0.20$), and Stroop ($F_{1,4894.41}=521.31$, $p<0.001$, $\eta^{2}_{p}=0.08$). Likewise, with the exception of N-Back 2 vs. 3 (t6114.86=-2.18, p=0.128, $\eta^{2}_{p}$=0.0), the time required to solve each task increased significantly as difficulty increased. See Table 3 for full post-hoc contrast results. In-person informal conversations with many participants indicated that they were feeling extremely challenged by 3-back such that even ‘giving up’ during the task at parts was common; to this end, the finding of similar reaction times during 2-back and 3-back is sensible.
Table 3: Post-hoc contrast results for modeling reaction time as a function of workload level for each task. Separate models were created for each task type; p-values are corrected using the Benjamini-Hochberg procedure. Outside of levels 2-3 in the N-Back task, reaction time increased significantly across difficulty levels of all tasks.
Task Contrast Est. SE df t p p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
Chess 0 - 3 -0.34 0.01 2913.17 -27.70 $<$0.001 $<$0.001 *** 0.21 [0.19,1]
Chess 0 - 2 -0.23 0.01 2886.14 -20.14 $<$0.001 $<$0.001 *** 0.12 [0.11,1]
Chess 1 - 3 -0.23 0.01 2911.83 -19.71 $<$0.001 $<$0.001 *** 0.12 [0.10,1]
Chess 0 - 1 -0.10 0.01 2868.08 -8.26 $<$0.001 $<$0.001 *** 0.02 [0.02,1]
Chess 1 - 2 -0.13 0.01 2909.87 -11.55 $<$0.001 $<$0.001 *** 0.04 [0.03,1]
Chess 2 - 3 -0.10 0.01 2821.98 -9.82 $<$0.001 $<$0.001 *** 0.03 [0.02,1]
N-Back 0 - 3 -0.17 0.00 6118.84 -41.79 $<$0.001 $<$0.001 *** 0.22 [0.21,1]
N-Back 0 - 2 -0.16 0.00 6124.60 -40.03 $<$0.001 $<$0.001 *** 0.21 [0.19,1]
N-Back 1 - 3 -0.10 0.00 6111.06 -24.76 $<$0.001 $<$0.001 *** 0.09 [0.08,1]
N-Back 0 - 1 -0.07 0.00 6115.00 -16.82 $<$0.001 $<$0.001 *** 0.04 [0.04,1]
N-Back 1 - 2 -0.09 0.00 6117.60 -22.74 $<$0.001 $<$0.001 *** 0.08 [0.07,1]
N-Back 2 - 3 -0.01 0.00 6114.86 -2.18 0.128 0.128 ns 0.00 [0.00,1]
Stroop 0 - 1 -0.07 0.00 4894.41 -22.83 $<$0.001 $<$0.001 *** 0.10 [0.08,1]
Rotation 0 - 3 -0.23 0.01 1552.22 -18.09 $<$0.001 $<$0.001 *** 0.17 [0.15,1]
Rotation 0 - 2 -0.19 0.01 1557.22 -15.37 $<$0.001 $<$0.001 *** 0.13 [0.11,1]
Rotation 1 - 3 -0.11 0.01 1558.35 -8.51 $<$0.001 $<$0.001 *** 0.04 [0.03,1]
Rotation 0 - 1 -0.12 0.01 1548.43 -9.85 $<$0.001 $<$0.001 *** 0.06 [0.04,1]
Rotation 1 - 2 -0.07 0.01 1548.64 -5.70 $<$0.001 $<$0.001 *** 0.02 [0.01,1]
Rotation 2 - 3 -0.04 0.01 1562.05 -2.90 0.020 0.021 * 0.01 [0.00,1]

### v2

@@ III-E Participant Information and Chess Player Skill
Chess participants also filled out a form indicating the frequency of their Chess play. The distribution of responses is shown in LABEL:supp-fig:PLAY_FREQUENCY in the supplementary material - a substantial majority (75%) of participants responded that they play at least once per week. After gathering our data, we assessed skill distribution of Chess players by examining the maximum puzzle difficulty each participant successfully completed, as illustrated in LABEL:supp-fig:MAX_ELO in the supplementary material. For reference, puzzles rated at 1400 Elo are the most frequently attempted puzzles in the lichess.org database [40]. While participant skill level could theoretically influence neural responses during puzzle solving, our preliminary analyses found no significant effects. Given these results, subsequent analyses do not investigate skill-based effects. Future research examining skill-based neural responses may benefit from recruiting a larger participant pool with broader representation across skill levels.
III-F EEG Data Collection and Preprocessing
@@ III-F1 Epoching the Data
III-F2 Frequency Domain Transformation and Prefrontal Averaging
Each epoch was transformed into the frequency domain with mne’s Multitaper library [51, 52]. Following the work of [11], we delineated frequency bands as follows: $\theta$ (4, 8), $\alpha_{1}$ (8, 11), $\alpha_{2}$ (11, 14), $\beta_{1}$ (14, 25), $\beta_{2}$ (25, 35), $\gamma_{1}$ (35, 40), and $\gamma_{2}$ (40, 45)555 While we recognize that these are not necessarily the most common waveband delineations used, they represent a similar paradigm within this work done in [11].. The Simpson integral [53] was used to extract single values representing the total power in each frequency band. Given our focus on overall pre-frontal activation patterns, we averaged the frequency-domain data from both probes into a single value per frequency bin, optimizing signal-to-noise ratio while maintaining measurement validity for our research objectives. Resulting power values were log-transformed.
III-F3 Data Collection Challenges and Solutions
Throughout our study we experienced connectivity issues with the MUSE device. In our initial implementation of the experimental protocol, this would cause complete data loss for a participant. To mitigate data loss, we implemented a reconnection protocol prior to the start of each task block (e.g., one block of N-Back, Stroop, mental workload, or Chess puzzles). This approach significantly reduced data loss overall, but some blocks were still lost for some participants. LABEL:supp-tab:DATA_DESCRIPTION in the supplementary material presents detailed information on the number of captured trials and epochs captured per-participant per-task after data loss was taken into account.
@@ III-G Workload Labels
III-H Statistical Methodology
The initial part of our analysis employs statistical methods to examine the measured neurophysiological data for both within-task workload and between-tasks. Except where otherwise stated, we utilize Linear Mixed-Effects models to achieve this aim. In each model, we incorporate Participant ID and Task Block per-participant as grouping factors for random effects. We implemented the models with R’s lmerTest package [54], and validated them through diagnostic checks that included verification of approximate normality of residuals with Quantile-Quantile (Q-Q) plots, posterior predictive checks for model fit, and examination of homoscedasticity and linearity against fitted values using the performance package [55]. To quantify the magnitude of the fixed effects, we calculate partial eta-squared ($\eta_{p}^{2}$) values using the effectsize library [56]. Estimated marginal means and post hoc pairwise comparisons were conducted with the emmeans library [57]. We used the conventional threshold of $\alpha=0.05$ for all statistical tests in our analysis.
III-I Modeling of Reaction Time and Correctness
To investigate behavioral metrics we modeled reaction time as a function of difficulty for all of the trials over all participants. For the N-Back, Stroop, and Rotation tasks, time is defined as the reaction time taken to press a key following the stimulus presentation. For the Chess puzzle task, we define time as the time required to solve the puzzle. Reaction time data was log-transformed for statistical modeling [58]. Omnibus test p-values for per-task models were corrected with Benjamini-Hochberg correction, as were all post-hoc comparisons. Prior to correction, pairwise comparisons were Tukey-adjusted to control the family-wise error rate. We also modeled correctness as a function of difficulty. Because correctness of a given trial is binary, we employed a Generalized Linear Mixed-Effect Model (glmer) with a binomial family function for the analysis; random effects structure is the same as described in Section III-H; p-values were adjusted using the same method as for the Reaction Time tests. Data visualizations are created using the seaborn library [59], and display group means accompanied by their 95% confidence intervals, which were calculated using bootstrapping methods that account for the hierarchical structure of the data.
III-J Statistical Modeling of EEG Data
@@
III-K Machine Learning Analyses
@@
III-K1 Grouped and Subsampled Monte Carlo Cross-Validation
Due to data loss from the MUSE device, the fact that the Chess task was collected for as long as the other three tasks combined, and due to varying time-spans of the other three tasks, our dataset contains imbalanced samples both within and across participants. However, this imbalance is not representative of the brain data itself; it stems solely from the experimental paradigm. To develop an online algorithm for differentiating between brain-states or workload levels in practice, we would modify the protocol to create a balanced dataset. Therefore, to test our data with machine learning in the fairest way possible we employed a grouped and subsampled Monte Carlo Cross-Validation methodology for both the within-task workload machine learning models and for the between-task model. The data preparation pipelines are shown in detail in the supplementary material in LABEL:supp-fig:ML_PIPELINE_WITHIN_TASK_WORKLOAD and LABEL:supp-fig:ML_PIPELINE_CROSS_TASK. This method involves 1,000 iterations for the cross-task workload model, and 1,000666For all of our models, after conducting 1,000 iterations of Monte Carlo Cross-Validation we determined that the 95 confidence interval of the mean fell within ±0.5 percentage point., where each iteration follows these steps:
@@ III-K1 Grouped and Subsampled Monte Carlo Cross-Validation
    Initialize a Random Forest classifier using the ensemble.RandomForest function from scikit-learn [60] on the training data using default parameters  [^7] Model configuration included 100 trees (n__\__estimators), unlimited maximum depth with nodes expanded until reaching pure leaf or minimum split threshold, minimum samples for split of 2, minimum samples per leaf of 1, and Gini criterion for measuring split quality. We opted for default parameters to establish a robust baseline implementation while avoiding potential overfitting from extensive parameter tuning; record both overall results as well as results of per-participant classification.
III-K2 Within-Task Workload Machine Learning
@@
For within-task workload we model each task separately. Because in this case we do not need to classify across tasks, we can leverage the entire dataset, including all valid data from participants in both phases I and II. See LABEL:supp-tab:PPM_WORKLOAD in the supplementary material for detailed information on average train and test set sizes across the Monte Carlo iterations.
III-K3 Cross-Task Machine Learning
@@
For cross-task analysis we require data for each participant from all tasks, therefore we use only the data collected in phase II for this machine learning set. For this machine learning set, we had a total of 6 participants. Across all Monte Carlo iterations we had an average of 2,723 training samples and 776 testing samples. Due to the data splitting process discussed in III-K1, the training and testing samples were each evenly balanced within each participant for samples per label, and were likewise balanced for samples per-participant. scikit-learn [60] was used to implement the grouping strategy, Random Forest model, and correctness statistics. imblearn [61] was used for subsampling of the data. We report the averages over all 1,000 models of the precision, recall, and macro F1-score.
IV Results
IV-A Reaction Time and Correctness
@@
See LABEL:supp-fig:BEHAVIORAL_STATS for a visualization of reaction time and correctness data. Regarding temporal statistics, for all tasks, time to solve was significant: N-Back ($F_{3,6166.49}=795.91$, $p<0.001$, $\eta^{2}_{p}=0.28$) [^8] All p-values reported inline reflect adjustments for multiple comparisons. Tables present both unadjusted (p) and adjusted (p.adj) values, Chess ($F_{3,2906.35}=300.37$, $p<0.001$, $\eta^{2}_{p}=0.24$), Rotation ($F_{3,1554.14}=127.65$, $p<0.001$, $\eta^{2}_{p}=0.20$), and Stroop ($F_{1,4894.41}=521.31$, $p<0.001$, $\eta^{2}_{p}=0.08$). Likewise, with the exception of N-Back 2 vs. 3 (t6114.86=-2.18, p=0.128, $\eta^{2}_{p}$=0.0), the time required to solve each task increased significantly as difficulty increased. See LABEL:supp-tab:BEHAVIORAL_TIME_POST_HOC in the supplementary material for full post-hoc contrast results. In-person informal conversations with many participants indicated that they were feeling extremely challenged by 3-back such that even ‘giving up’ during the task at parts was common; to this end, the finding of similar reaction times during 2-back and 3-back is sensible.